OpenStudy (anonymous):

Proof that x^y + y^x > 1 for all x,y > 0

OpenStudy (anonymous):

x,y>0 so , x^y>0 and y^x>0 both of them cannot be 0 and so it is x^y + y^x > 1 hence proved

OpenStudy (asnaseer):

@Dhashi - how do you conclude x^y+y^x>1 if each component is greater than 0?

OpenStudy (jamesj):

that's not a proof.

OpenStudy (anonymous):

why can't they both be 1/4?

OpenStudy (anonymous):

It was a nice try though ;)

OpenStudy (aravindg):

positive quantity raised to positive quantity always positive

OpenStudy (anonymous):

@jamesj, for some reason i feel like lighting up a marlboro

OpenStudy (anonymous):

AravindG and DHASHNI propably didnt see that the RHS was 1, not 0

OpenStudy (asnaseer):

It looks like a very interesting problem - I wonder if there is a simple way of proving it (apart from the one @satellite73 showed in the link)?

OpenStudy (asnaseer):

just wondering if we could use logs some how?

OpenStudy (jamesj):

@sat73, Phillip Morris paid me to use that photo. I can't seem to be able to read the proof you linked to, and it's being difficult about registering. Would you mind taking a quick screen shot of the text and posting it?

OpenStudy (anonymous):

James use your mouse.

OpenStudy (anonymous):

Hahaha asnaseer quite a good result using logs hehehe if x and y are 1 tehn using log we will remain with 0>0 and the question itself may be wrong it might be > or equal to 1

OpenStudy (jamesj):

ah, select the text, got it.

OpenStudy (anonymous):

mouse over, or you can look here i think it is ok http://www.ocf.berkeley.edu/~wwu/cgi-bin/yabb/YaBB.cgi?board=riddles_medium;action=print;num=1213001311

OpenStudy (anonymous):

hmmmm sry sry

OpenStudy (asnaseer):

This is the text from that site: Assume that both x and y lie in (0,1) (otherwise, we are done). Now wlog assume that y1/(1-y) >= x1/(1-x) Note that, this implies that xy-1/yx-1 >= 1 -- ( A ) Now Consider S = (xy + yx)1/x = [yx(1 + xy/yx)]1/x = y(1 + xy/yx)1/x Now since 1/x > 1, by Bernoulli's inequality that (1+x)r >= 1 + rx for x > -1 and r >= 1 we have that S >= y(1 + xy-1/yx) = y + xy-1/yx-1 >= y + 1 (by A) Thus S >= 1 + y > 1 Since S > 1, Sx > 1 and we are done.

OpenStudy (mr.math):

This is also from the same site: Case 1: Either x,y >= 1 then x^y >=1 or y^x >=1 => x^y + y^x >1 Case 2: x,y<1 By Bernoulli inequality, we have: x^y = [1 + (x-1)]^y >= 1 + y(x-1) => x^y + y -1 >= xy Likewise: y^x + x-1 >= xy Thus, if we can show 2xy+1>= x+y the x^y + y^x > 1. This is true since (x-1)(y-1)+xy>0 given <0x,y<1

OpenStudy (anonymous):

Got one thing wrong with logs i have assumed wrong sry guys there's no prob with the question

OpenStudy (anonymous):

Y cant we guys just differentiate and find the minimum value of the function???????

OpenStudy (anonymous):

Using some Tyler expansion or that kinda thing for the function of two variables

OpenStudy (anonymous):

Asnaseer I need ur comment here

OpenStudy (anonymous):

I tried differentating, but its rather messy, so I didnt want to go on that. The question is something my friend asked me on facebook couple of weeks ago, and I couldnt solve.

OpenStudy (aravindg):

pls help me in physics after this guys.come to physics section????

OpenStudy (asnaseer):

@Atchyut - what do you want me to comment on?

OpenStudy (anonymous):

@asnaseer: I didn't understand: Now wlog assume that y1/(1-y) >= x1/(1-x) what's is wlog?

OpenStudy (asnaseer):

hmmm - I just copied the text from the site that @satellite73 gave above.

OpenStudy (anonymous):

I know, but you can you use your amazing mind to figure it out though ;)

OpenStudy (jamesj):

wlog means without loss of generality

OpenStudy (asnaseer):

thx @JamesJ

OpenStudy (anonymous):

argghhh that's silly abbreviation, I was thinking some sort of mathematical function or something.